OpenStudy (peterg2001):

Which of the following expressions is equivalent to 6m - 10? 2(3m + 5) -2(5 - 3m) 2(3m - 10)

OpenStudy (peterg2001):

@Monkey4224

OpenStudy (anonymous):

2(3m-5)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Take each choice and multiply it out using the distributive property. A. 2(3m + 5) = 6m + 10 (not correct) B. -2(5 - 3m) = -10 + 6m = 6m - 10 (looking good) C. 2(3m - 10) = 6m - 20 (not correct) B. is correct .